OpenStudy (hahawowoh):

Will fan and medal! Write the equation of the line that passes through the points (3, 6) and (5, 18) using function notation. f(x) = 2x + 12 f(x) = 6x − 12 y = 2x + 12 y = 6x − 12

Mehek (mehek14):

function notation = \(\tt{a(b)}\) so not C or D

Mehek (mehek14):

\(\tt{\dfrac{18-6}{5-3}=\dfrac{12}{2}=6}\)

Mehek (mehek14):

so B

OpenStudy (hahawowoh):

thank you so much!!

Mehek (mehek14):

yw:)